# LinearSR: Unlocking Linear Attention for Stable and Efficient Image Super-Resolution

### v2

@@ Abstract
Generative models for Image Super-Resolution (SR) are increasingly powerful, yet their reliance on self-attention’s quadratic complexity ($O(N^{2})$) creates a major computational bottleneck. Linear Attention offers an $O(N)$ solution, but its promise for photorealistic SR has remained largely untapped, historically hindered by a cascade of interrelated and previously unsolved challenges. This paper introduces LinearSR, a holistic framework that, for the first time, systematically overcomes these critical hurdles. Specifically, we resolve a fundamental, training instability that causes catastrophic model divergence using our novel “knee point”-based Early-Stopping Guided Fine-tuning (ESGF) strategy. Furthermore, we mitigate the classic perception-distortion trade-off with a dedicated SNR-based Mixture of Experts (MoE) architecture. Finally, we establish an effective and lightweight guidance paradigm, TAG, derived from our “precision-over-volume” principle. Our resulting LinearSR model simultaneously delivers state-of-the-art perceptual quality with exceptional efficiency. Its core diffusion forward pass (1-NFE) achieves SOTA-level speed, while its overall multi-step inference time remains highly competitive. This work provides the first robust methodology for applying Linear Attention in the photorealistic SR domain, establishing a foundational paradigm for future research in efficient generative super-resolution.
@@ 1 Introduction
Recent advancements in Image Super-Resolution (SR) are dominated by generative models (Chen et al., 2023; Rombach et al., 2022) that leverage the powerful self-attention mechanism to synthesize photorealistic details. However, this power comes at a steep price: the quadratic complexity ($O(N^{2})$) of self-attention imposes a major computational bottleneck. Linear Attention (Shen et al., 2021; Katharopoulos et al., 2020; Wang et al., 2020; Cai et al., 2022), with its $O(N)$ complexity, has emerged as a compelling alternative. Its potential was successfully shown in general image generation by models like SANA (Xie et al., 2024), which validated its ability to capture global dependencies efficiently. This work addresses the central challenge: how can the efficiency of Linear Attention be fully unlocked to satisfy the extreme fidelity requirements of super-resolution?
@@ 4.1 Experimental Settings
For the pre-training and MoE stages, we use the public datasets DIV2K (Agustsson & Timofte, 2017), LSDIR (Li et al., 2023), and ReLAION-High-Resolution (laion, 2024), along with a custom high-resolution dataset crawled from Unsplash. The final SFT stage employs a curated set of high-quality images collected from the Internet. We synthesize 4x LR-HR pairs (256×256 from 1024×1024) using the Real-ESRGAN (Wang et al., 2021) degradation pipeline. Our model is trained with a flow-based objective (Lipman et al., 2022), employing the CAME optimizer (Luo et al., 2023) with a constant learning rate of 1e-4 and a batch size of 14. Stage 1 utilized 8 NVIDIA A800 GPUs, while each of the four experts in Stage 2 was trained on 6 A800 GPUs. Following our ESGF strategy, each stage proceeds until its performance ‘knee-point’ is reached. For inference, we use 20 sampling steps to generate results.
@@
We conduct evaluations on RealSR (Cai et al., 2019), DrealSR (Wei et al., 2020), RealLQ250 (Ai et al., 2024), and a synthetic test set of 100 images from DIV2K-Val (Agustsson & Timofte, 2017) generated with our training degradation. For RealSR and DrealSR, we follow the established protocol (Wu et al., 2024a; b; Chen et al., 2025a) of center-cropping inputs to 256×256. We compare LinearSR against ten state-of-the-art methods: StableSR (Wang et al., 2024a), DiffBIR (Lin et al., 2024), SeeSR (Wu et al., 2024b), SUPIR (Yu et al., 2024), DreamClear (Ai et al., 2024), SinSR (Wang et al., 2024b), OSEDiff (Wu et al., 2024a), AdcSR (Chen et al., 2025a), InvSR (Yue et al., 2025), and TSD-SR (Dong et al., 2025).
@@ 4.2 Comparison with State-of-the-Arts
1 Image Inference Time (s) 78.405 25.543 13.632 13.808 16.319 8.999 1.086 0.561 0.667 12.635 0.830
@@ 4.3.2 Necessity of the ESGF Strategy for Stable Fine-tuning
Baseline
N/A N/A 24.85 0.691 0.374 0.363 63.93 0.572
(b) 2-Expert MoE
SNR-based
[0.875] 25.02 0.671 0.377 0.374 63.18 0.591
(c)Ours
4-Expert MoE
SNR-based
[0.223, 0.875,0.939]
25.00 0.682 0.375 0.371 64.02 0.598
4-Expert MoE
Naive Uniform
[0.25, 0.5,0.75]
24.84 0.666 0.389 0.368 62.51 0.582
@@ Appendix C User Study
Appendix D More Visual Comparisons
@@
To further demonstrate the robustness and superiority of our proposed LinearSR, we provide extensive visual comparisons against state-of-the-art diffusion-based methods. Figures 8 and 9 showcase results on challenging real-world and synthetic degradation scenarios, respectively. These qualitative results comprehensively validate our method’s effectiveness.

### v4

@@ Abstract
Generative models for Image Super-Resolution (SR) are increasingly powerful, yet their reliance on self-attention’s quadratic complexity ($O(N^{2})$) creates a major computational bottleneck. Linear Attention offers an $O(N)$ solution, but its promise for photorealistic SR has remained largely untapped, historically hindered by a cascade of interrelated and previously unsolved challenges. This paper introduces LinearSR, a holistic framework that, for the first time, systematically overcomes these critical hurdles. Specifically, we resolve a fundamental, training instability that causes catastrophic model divergence using our novel “knee point”-based Early-Stopping Guided Fine-tuning (ESGF) strategy. Furthermore, we mitigate the classic perception-distortion trade-off with a dedicated SNR-based Mixture of Experts (MoE) architecture. Finally, we establish an effective and lightweight guidance paradigm, TAG, derived from our “precision-over-volume” principle. Our resulting LinearSR model simultaneously delivers state-of-the-art perceptual quality with exceptional efficiency. Its core diffusion forward pass (1-NFE) achieves SOTA-level speed, while its overall multi-step inference time remains highly competitive. This work provides the first robust methodology for applying Linear Attention in the photorealistic SR domain, establishing a foundational paradigm for future research in efficient generative super-resolution. Code and models are available: https://github.com/xh9998/LinearSR.
@@ 1 Introduction
Recent advancements in Image Super-Resolution (SR) are dominated by generative models (Chen et al., 2023; Rombach et al., 2022; Duan et al., 2025) that leverage the powerful self-attention mechanism to synthesize photorealistic details. However, this power comes at a steep price: the quadratic complexity ($O(N^{2})$) of self-attention imposes a major computational bottleneck. Linear Attention (Shen et al., 2021; Katharopoulos et al., 2020; Wang et al., 2020; Cai et al., 2022), with its $O(N)$ complexity, has emerged as a compelling alternative. Its potential was successfully shown in general image generation by models like SANA (Xie et al., 2024), which validated its ability to capture global dependencies efficiently. This work addresses the central challenge: how can the efficiency of Linear Attention be fully unlocked to satisfy the extreme fidelity requirements of super-resolution?
@@ 4.1 Experimental Settings
For the pre-training and MoE stages, we use the public datasets DIV2K (Agustsson and Timofte, 2017), LSDIR (Li et al., 2023), and ReLAION-High-Resolution (laion, 2024), along with a custom high-resolution dataset crawled from Unsplash. The final SFT stage employs a curated set of high-quality images collected from the Internet. We synthesize 4x LR-HR pairs (256×256 from 1024×1024) using the Real-ESRGAN (Wang et al., 2021) degradation pipeline. Our model is trained with a flow-based objective (Lipman et al., 2022), employing the CAME optimizer (Luo et al., 2023) with a constant learning rate of 1e-4 and a batch size of 14. Stage 1 utilized 8 NVIDIA A800 GPUs, while each of the four experts in Stage 2 was trained on 6 A800 GPUs. Following our ESGF strategy, each stage proceeds until its performance ‘knee-point’ is reached. For inference, we use 20 sampling steps to generate results.
@@
We conduct evaluations on RealSR (Cai et al., 2019), DrealSR (Wei et al., 2020), RealLQ250 (Ai et al., 2024), and a synthetic test set of 100 images from DIV2K-Val (Agustsson and Timofte, 2017) generated with our training degradation. For RealSR and DrealSR, we follow the established protocol (Wu et al., 2024a; b; Chen et al., 2025a) of center-cropping inputs to 256×256. We compare LinearSR against ten state-of-the-art methods: StableSR (Wang et al., 2024a), DiffBIR (Lin et al., 2024), SeeSR (Wu et al., 2024b), SUPIR (Yu et al., 2024), DreamClear (Ai et al., 2024), SinSR (Wang et al., 2024b), OSEDiff (Wu et al., 2024a), AdcSR (Chen et al., 2025a), InvSR (Yue et al., 2025), and TSD-SR (Dong et al., 2025).
@@ 4.2 Comparison with State-of-the-Arts
1 Image Inference Time (s) 78.405 25.543 13.632 13.632 16.319 8.999 1.086 0.561 0.667 12.635 0.830
@@ 4.3.2 Necessity of the ESGF Strategy for Stable Fine-tuning
(a) Baseline N/A N/A 24.85 0.691 0.374 0.363 63.93 0.572
(b) 2-Expert MoE SNR-based [0.875] 25.02 0.671 0.377 0.374 63.18 0.591
(c)Ours 4-Expert MoE SNR-based [0.223, 0.875,0.939] 25.00 0.682 0.375 0.371 64.02 0.598
(d) 4-Expert MoE Naive Uniform [0.25, 0.5,0.75] 24.84 0.666 0.389 0.368 62.51 0.582
@@ 5 Conclusion
Acknowledgments
This work is supported by Shanghai Artificial Intelligence Laboratory.
Ethics Statement
Our LinearSR is the first framework to successfully unlock the potential of linear attention for high-fidelity super-resolution. To ensure ethical compliance, our training data was curated from a combination of established public datasets and publicly accessible online platforms, such as Unsplash. We have taken deliberate measures to minimize potential biases in this data, fully aligning with universal ethical guidelines. We explicitly emphasize that this framework is not intended for misuse in achieving harmful purposes; downstream users are encouraged to adhere to ethical principles when applying the technology. Additionally, all authors declare no conflicts of interest related to this work.
Reproducibility Statement
To enable direct reproduction of the quantitative metrics in our tables and the visual comparisons in our figures, we plan to release the corresponding pre-trained model checkpoints and evaluation scripts, pending institutional approval. To ensure a consistent testing environment, we will also provide detailed specifications of the hardware used, such as GPU models, and the exact versions of all software libraries. This commitment ensures that our reported results can be precisely verified by the community.
@@ Appendix C User Study
Appendix D Robustness of the Knee-Point Selection Strategy
In this section, we elaborate on the rigorous methodology used to determine the optimal training termination step, referred to as the “Knee-Point.” Rather than relying on arbitrary heuristics or manual guesswork, we employ a systematic two-stage strategy: (1) Automated Detection via Metric Variance Analysis, followed by (2) Human Verification.
D.1 Automated Detection Algorithm
To identify the precise transition from the convergence phase to the oscillation phase, we track 6 metrics (PSNR, SSIM, LPIPS, MANIQA, MUSIQ, CLIPIQA) across 4 diverse datasets (DIV2K-Val, RealSR, DrealSR, RealLQ250). This covers both synthetic and real-world distributions.
We define the Knee-Point $t^{*}$ as the critical step where the model achieves maximum stability before performance degradation begins. The detection process is formalized in Algorithm 1. Let $\mathcal{M}=\{m_{1},m_{2},\dots,m_{T}\}$ be the sequence of a validation metric recorded at steps $t$. We utilize a sliding window of size $W$ to monitor the local variance and the future trend slope.
Algorithm 1 Automated Knee-Point Detection Strategy
0: Metric sequence $\mathcal{M}=\{m_{t}\}_{t=1}^{T}$, Window size $W$, Stability threshold $\epsilon_{stable}$
0: Optimal Knee-Point $t^{*}$
1: Initialize candidates set $\mathcal{C}\leftarrow\emptyset$
2:for $t=W$ to TWT-W do
3:  {Calculate local variance over the past window}
4:  $V_{t}\leftarrow\text{Var}(m_{t-W:t})$
5:  {Calculate slope trend over the future window}
6:  $S_{t}\leftarrow\text{Slope}(m_{t:t+W})$
7:  {Identify stability region before negative trend}
8:  if $V_{t}<\epsilon_{stable}$ and $S_{t}<0$ then
9:   $\mathcal{C}\leftarrow\mathcal{C}\cup\{t\}$
10:  end if
11:end for
12: {Select the latest step satisfying criteria}
[…]
14:return $t^{*}$
Following the automated proposal of $t^{*}$, we conduct a Human Verification phase. We visually inspect the validation curves and generated samples around $t^{*}$ to ensure the detected point does not correspond to a temporary local fluctuation but represents a genuine structural convergence.
D.2 Empirical Validation and Universality
To demonstrate the universality of this phenomenon, we present the training log of a separate optimization experiment in Table 8 and visualize the trend in Figure 8.
While the Knee-Point in our main paper was detected at 48k steps, this separate experimental run exhibits the Knee-Point at 58k steps. Despite the shift in the absolute step number due to different hyperparameter settings, the underlying pattern remains consistent:
  1. 1.
    Rapid Ascent Phase: Metrics improve significantly in the early stages.
  2. 2.
    Knee-Point ($t^{*}$): Performance peaks and variance is minimized (highlighted in bold in Table 8).
  3. 3.
    Oscillation Phase: Beyond $t^{*}$, the metrics enter a phase of high variance or slight degradation, indicating overfitting or instability inherent to the adversarial/perceptual loss components.
As shown in Table 8, at step 58k, the model achieves the optimal trade-off across perception (MANIQA, MUSIQ) and fidelity (PSNR, SSIM) metrics. For instance, on the RealSR dataset, MUSIQ peaks at 60.231 at 58k before dropping to 51.643 at 242k, confirming the necessity of our early stopping strategy.
Figure 8: Visualization of the Knee-Point Phenomenon. The curves illustrate the evolution of 6metrics throughout the training process. A distinct “Knee-Point” is observable around step 58k, marked by a peak in performance and stability, followed by an oscillation phase where metrics fluctuate or degrade. This visual evidence corroborates the quantitative data in Table 8.
Appendix E Benchmarking Protocols and Quality-Latency Alignment
In this section, we provide a detailed clarification of our benchmarking methodology, specifically focusing on the measurement of inference latency and the structural differences between the original and distilled models.
E.1 Definition of 1-NFE Forward Time
The reported “1-NFE Forward Time” in our experiments strictly measures the core diffusion denoising step. This measurement excludes auxiliary computational costs such as the Text Encoder, VAE decoding, and pre-processing steps. This isolation ensures that the metric purely reflects the architectural efficiency of the diffusion backbone itself.
E.2 The “Quality-Latency Alignment” Standard
To ensure a rigorous and fair comparison, we adhere to a Quality-Latency Alignment standard. This principle dictates that the latency reported for a model must be measured using the exact configuration (e.g., hyperparameters, guidance strategies) required to achieve the optimal visual quality presented in the qualitative comparisons.
Impact of Classifier-Free Guidance (CFG).
A critical factor influencing this measurement is the use of Classifier-Free Guidance (CFG). Based on this principle, the “unit work” per step differs across methods:
  • Methods with CFG (Heavier): Like LinearSR, DiffBIR, DreamClear, SeeSR, and SUPIR, these models default to using CFG. Their “core denoising step” physically necessitates computing both conditional and unconditional branches (effective batch size = 2).
  • Methods without CFG (Lighter): Distilled or single-branch methods like AdcSR, OSEDiff, InvSR, TSD-SR, and SinSR typically use a single branch (effective batch size = 1).
E.3 Efficiency Analysis
It is acknowledged that single-branch models naturally possess a computational advantage in 1-NFE time due to their reduced workload. Despite this structural difference, our model achieves a state-of-the-art 1-NFE efficiency of 0.036s. This performance highlights the intrinsic speed of LinearSR. It indicates that our architectural optimizations are orthogonal to distillation, suggesting significant potential for further acceleration when combined with such techniques.
Appendix F More Visual Comparisons
@@
To further demonstrate the robustness and superiority of our proposed LinearSR, we provide extensive visual comparisons against state-of-the-art diffusion-based methods. Figures 9 and 10 showcase results on challenging real-world and synthetic degradation scenarios, respectively. These qualitative results comprehensively validate our method’s effectiveness.
@@
Table 8: Full training log across all steps. The 58k step is highlighted as the Knee-Point.
Datasets Metrics 4k 12k 20k 28k 36k 44k 50k 58k 66k 70k 74k 80k 82k 96k 128k 142k 158k 180k 196k 212k 226k 242k
[…]
@@
Appendix G Declaration of Use of Large Language Models (LLM)
We affirm that this paper was primarily written by the authors. Large Language Models (LLMs) were utilized solely as general-purpose assistive tools for language refinement, grammar correction, and stylistic improvements during the writing process. Specifically, Gemini 2.5 Pro (Comanici et al., 2025) was employed for minor text polishing and rephrasing to enhance clarity and readability. No LLM was used for conceptual ideation, experimental design, data analysis, or generating any substantive content of the research.